OpenStudy (anonymous):

whats 0.33, 4/7, and 1/7 In order from greatest to least

OpenStudy (lizz123):

divide 7 by 4 to get it to a decimal.

OpenStudy (lizz123):

So 4/7 is 1.75 in a decimal

OpenStudy (lizz123):

Now figure out 1/7 into a decimal form

OpenStudy (lizz123):

7/1 into a decimal form is 7

OpenStudy (lizz123):

So now you have 0.33 , 1.75 , 7

OpenStudy (lizz123):

Is 0.33 bigger than 1.75? No it is not so 0.33 goes first. Is 1.75 bigger than 7? No it is not so 4/7 goes next. Now you are left with 7 so you have 1/7. The order goes: \[0.33, \frac{ 4 }{ 7 }, \frac{ 1 }{ 7 }\]

OpenStudy (lizz123):

I mean the order would be : \[\frac{ 1 }{ 7 }, \frac{ 4 }{ 7 }, 0.33\]

OpenStudy (lizz123):

Do you still need help